Please help me solve this: I have tried numerous differnt ways and keep getting nowhere. The problem is : Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
-5x + 4y = -16
-15x + 2y = -58

Answer by jim_thompson5910(35256)

Hint: Multiply EVERY term in the second equation by -2 to go from

-15x + 2y = -58

to

30x - 4y = 116

-------------------------------------------------------

So

-5x + 4y = -16
-15x + 2y = -58

turns into

-5x + 4y = -16
30x - 4y = 116